OpenStudy (anonymous):

Is the line through points P(4,9) and Q(1,8) parallel to the line through points R(-1,2) and S(-5,0)? Explain A) no; one line has zero slope, the other has no slope B) yes; the lines have equal slops C) yes; the lines are both vertical D) no; the lines have unequal slopes

OpenStudy (anonymous):

Slope of PQ = (8-9)/(1-4) = 1/3 Slope of RS = (0-2)/(-5 - -1) = -2/-4 = 1/2 Since the slopes are UNEQUAL, the lines are NOT parallel. Choice (D).